Let E be the event that a corn crop has an infestation of ear worms, and let B be the event that a corn crop has an infestation of corn borers.

Suppose that P(E) = 0.29, P(B) = 0.1, and P(E and B) = 0.09. Find the probability that a corn crop has no corn borer infestation.

The answer key says that the answer is 0.9 and I cannot figure out why. I would appreciate any help or explanation, thanks.
Answer #1

It's given P(B) = 0.1

B --> event of corn borers.

P(no corn borers) = 1 - P(corn borers) = 1 - P(B) = 1 - 0.1 = 0.9

Hence, 0.9 is the probability that a corn crop has no corn borer infection.

Please comment if any doubt. Thank you.